If you drove to Austin at 70 miles per hour and returned at 50 miles per hour.what is the distance driven to Austin if total driving time was 7.2 hours
Found 2 solutions by nshah11, Alan3354:
Answer by nshah11(47)

Assume the distance to Austin is d miles.
d*(1/70 + 1/50) = 36/5
d = (36/5)*(50*70/120) = 210 miles

Answer by Alan3354(69443)

If you drove to Austin at 70 miles per hour and returned at 50 miles per hour.what is the distance driven to Austin if total driving time was 7.2 hours
-----------
Avg speed for the round trip = 2*70*50/(70+50)
= 7000/120
= 175/3 mi/hr
------------
d = r*t = (175/3)*7.2
d = 420 miles RT
one-way = 210 miles